# Mapping the Potential of Explainable Artificial Intelligence (XAI) for Fairness Along the AI Lifecycle

### v2

@@ 3. How XAI Can Be Leveraged for Fairness Along the AI Lifecycle
Grounded in a broad interdisciplinary literature review (summarized in Table 1), we distill eight categories of what previous work on fairness has called for. We call these categories fairness desiderata, and each desideratum can be instantiated with specific fairness objectives. For the purposes of this paper, we focus specifically on fairness desiderata connected to XAI. While there are various stakeholders involved in utilizing XAI for fairness, an extensive discussion of their various roles is beyond the scope of this paper. When we mention stakeholders, we rely on the taxonomy provided by Langer et al. (2021b).
@@
Table 1. Fairness desiderata and their related concepts in existing interdisciplinary literature.
@@
Understanding bias (Ntoutsi et al., 2020)
Interdisciplinary fairness conceptualization (Mulligan et al., 2019)
Lessons from political philosophy (Binns, 2018)
Socio-technical perspective (Sartori and Theodorou, 2022)
@@
Sampling bias (Mehrabi et al., 2021; De-Arteaga et al., 2022)
Data errors and bias (Pradhan et al., 2022)
Data-centric factors in algorithmic fairness (Li et al., 2022)
@@
Algorithmic bias (De-Arteaga et al., 2022)
Formal fairness definitions (Corbett-Davies and Goel, 2018)
Fairness metrics (Castelnovo et al., 2022)
Statistical fairness criteria (Barocas et al., 2019)
Disparate impact & disparate treatment (Barocas and Selbst, 2016)
@@
Fairness perceptions (Starke et al., 2022a; Schoeffer et al., 2022)
Perceptions of justice (Colquitt, 2001; Binns et al., 2018)
Fairness judgment (Dodge et al., 2019)
Society-in-the loop (Rahwan, 2018)
@@
Human-ML augmentation for fairness (Teodorescu et al., 2021)
Appropriate reliance (Schemmer et al., 2023)
Human-in-the-loop (Dodge et al., 2019)
Domain expertise (Chakraborty et al., 2020)
@@
Self-informed advocacy (Vredenburgh, 2022)
Procedural fairness (European Commission. Directorate General for Communications Networks, Content and Technology. and High Level Expert Group on Artificial Intelligence., 2019)
Counterfactual explanations (Wachter et al., 2017)
Fair and adequate explanations (Asher et al., 2022)
The explanation game (Watson and Floridi, 2021)
@@
Long-term effects of algorithmic fairness (De-Arteaga et al., 2022)
Fairness drift (Ghosh et al., 2022)
Fairness through time (Castelnovo et al., 2021)
Fairness monitoring (Hardt et al., 2021; Panigutti et al., 2021)
@@
Informational fairness (Colquitt, 2001)
Design publicity (Loi et al., 2021)
Model cards (Mitchell et al., 2019)
Outward transparency (Walmsley, 2021)
Figure 2. Fairness desiderata along the AI lifecycle depicting how XAI may directly contribute to these desiderata and how it may contribute to informational fairness across all desiderata (symbolized by speakerphones).
3.1. Fairness Understanding
“Accounting for bias not only requires an understanding of the different sources, that is, data, knowledge bases, and algorithms, but more importantly, it demands the interpretation and description of the meaning, potential side effects, provenance, and context of bias.” (Ntoutsi et al., 2020, p. 8)
@@ 3.1. Fairness Understanding
Among others, this may include a sound understanding of fairness schools of thought (Mulligan et al., 2019), existing social inequalities (Green, 2022), stakeholder-centered fairness requirements (Langer et al., 2021b), or relevant legal frameworks (Hacker and Passoth, 2022). As the following fairness desiderata will show, there is no one-size-fits-all way in which a system can be fair, both across (e.g., a system can be perceived as fair despite not being formally fair) as well as within fairness desiderata (e.g., formal fairness criteria pose inherent tradeoffs (Friedler et al., 2021)).
@@
The Role of XAI. XAI can contribute to a better understanding of fairness across all fairness desiderata which, e.g., may lead to a re-evaluation of earlier fairness objectives. Data-centric explanations (Anik and Bunt, 2021) might reveal pre-existing group disparities (e.g., in the form of differing base rates such as the gender pay gap), which are crucial factors in determining the fairness objectives and how to achieve them. Further, simulating and testing various formal fairness metrics promotes an understanding of conflicting objectives that may lead to adjustments or re-weightings of fairness objectives (De-Arteaga et al., 2022). Similarly, XAI may change stakeholders’ view of proactively using sensitive features in a specific decision-making context (Nyarko et al., 2021). With regard to “substantive” fairness (Green, 2022; Khan et al., 2022), XAI may also be applied to normatively gauge the legitimacy of certain features depending on the societal context.
3.2. Data Fairness
@@
“Explanations […] are crucial for helping system developers and ML practitioners to debug ML algorithms by identifying data errors and bias in training data, such as measurement errors and misclassifications, data imbalance, missing data and selection bias, covariate shift, technical biases introduced during data preparation, and poisonous data points injected through adversarial attacks.” (Pradhan et al., 2022, p. 248)
@@
Undesirable model behavior often stems from flawed data that contains misrepresentations of the world (e.g., erroneous, mislabeled, or imbalanced data) or accurate representations that are societally undesirable (e.g., historical inequalities). The goal of data fairness is to identify and address such flaws in the data set used for the training of an AI model.
@@
Prior studies have pointed out data issues as drivers of unfairness (Barocas et al., 2017; Mehrabi et al., 2021; Li et al., 2022) or highlighted data as a starting point for unfairness mitigation (Li and Vasconcelos, 2019; Cai et al., 2020; Saleiro et al., 2020). Mehrabi et al. (2021) provide a comprehensive list of data biases that may introduce unfairness early on in AI development. Awareness and understanding of these biases are crucial to derive strategies to handle them.
@@
Stage of the Lifecycle. Data fairness relates to data collection (e.g., in the form of labeling errors, sampling bias, imbalances, etc.) and data analysis, which aims to identify and potentially mitigate unfair data characteristics early on. However, data fairness also reaches into feature selection which is usually part of an iterative loop together with model construction. For example, features might be dropped if they are not justifiably task-relevant. Importantly, in the context of sensitive attributes, developers should always be aware of the flaws of the idea of “fairness through unawareness” (Dwork et al., 2012; Nyarko et al., 2021).
@@
The Role of XAI. As data is a primary source of unfairness, XAI is suited to identify potential disparities, imbalances, or abnormalities manifested in the available data early on. Descriptive statistics are a natural first step to explore pre-existing disparities (Li et al., 2022). Anik and Bunt (2021) and Mitchell et al. (2019) provide two examples of how to present simple descriptions and visualizations about the collection, feature distributions, and patterns of a dataset. XAI techniques have further been claimed to reveal instances and features in the data that have undesirable effects on the model output (Black et al., 2020; Fan et al., 2022; Pradhan et al., 2022) which, however, are often subject to questionable causal and normative assumptions (Chiappa, 2019; Binns, 2020). These approaches indicate a strong connection between data fairness and formal fairness.
3.3. Formal Fairness
@@
“To counteract biases, it is, therefore, crucial to enable their detection. Explainability approaches may aid in this regard by providing means to track down factors that may have contributed to unfair and unethical decision-making processes and either to eliminate such factors, to mitigate them, or at least to be aware of them.” (Langer et al., 2021b, p. 6)
@@
The most common fairness desideratum is concerned with formal model properties. By formal fairness, we refer to the vast array of formal criteria that have been proposed as mathematical and statistical measures of fairness (Barocas et al., 2019; Castelnovo et al., 2022).
@@
Consistent with Verma and Rubin (2018), this includes all fairness definitions based on statistical measures (e.g., demographic parity), similarity-measures (e.g., fairness through awareness), or based on causal reasoning (e.g., counterfactual fairness). Formal fairness notions are often distinguished into group and individual fairness. Group fairness criteria typically require a form of parity between demographic groups, e.g., along sensitive attributes like gender or race (Chouldechova, 2017). Individual fairness criteria typically demand to treat similar people alike (Dwork et al., 2012).
@@
Stage of the Lifecycle. Formal fairness is particularly relevant for the iterative loop of model construction and model evaluation. As soon as the first model prototype is ready, fairness metrics can be evaluated. Based on the evaluation, unfairness mitigation techniques can be implemented and validated iteratively (Chen et al., 2023).
@@
The Role of XAI. Regarding formal fairness, XAI could be of exploratory value offering a plethora of tools to explore formal fairness from multiple angles. Again, descriptive statistics present a natural first step to identify potential disparities, imbalances, or abnormalities outputted by the model (Chouldechova, 2017; Ahn and Lin, 2020). This is especially useful if formal fairness objectives are already specified, e.g., when testing whether formal group fairness metrics are sufficiently satisfied (Chen et al., 2023). XAI might also point towards specific features driving the violation of group fairness metrics (Ghosh et al., 2023) or shed light on more subtle forms of formal fairness such as fairness of recourse (Gupta et al., 2019). XAI has further been claimed to elucidate the complex interplay between “task-relevant” and correlated “protected” features (e.g., (Grabowicz et al., 2022)) which, again, relies on causal and normative assumptions (Chiappa, 2019; Binns, 2020). Technical possibilities of XAI to explore formal fairness are manifold but require utmost caution when applied for unfairness mitigation in specific application contexts and should not be misinterpreted as fairness “proofs” or “guarantees“ (Deck et al., 2023). Particularly, when interpreting the legitimacy of using sensitive information, it is paramount to account for the “fairness through unawareness” fallacy (Dwork et al., 2012; Lipton et al., 2018) and for differential subgroup validity (Hunter et al., 1979; Gupta et al., 2023).
3.4. Perceived Fairness
@@
“We argue that only fair systems that are also perceived to be fair by their users should be accepted and employed in practice.” (Shulner-Tal et al., 2022a, p. 5)
@@
Another common fairness desideratum refers to positive perceptions of stakeholders, which is often related to trust and acceptance (Shulner-Tal et al., 2022a; Papenmeier et al., 2019). In this sense, perceived fairness captures how stakeholders, particularly those affected by a decision, perceive the fairness of the AI system.
@@
Measures of perceived fairness can be derived from the justice constructs of Colquitt (2001), which decompose fairness perceptions into a procedural, distributive, and informational dimension (see also (Binns et al., 2018; Dodge et al., 2019; Schoeffer et al., 2022)). Accounting for fairness perceptions promotes a valuable means to design AI systems based on human needs and ideals while giving voice to societal values which Rahwan (2018) coined as “society-in-the-loop”. Notably, the desirable aspect of perceived fairness does not necessarily or solely lie in positive fairness perceptions but in “appropriate” fairness perceptions (Schoeffer and Kuehl, 2021).
@@
Stage of the Lifecycle. Although fairness perceptions can be relevant at any stage of the lifecycle (e.g., to evaluate data fairness (Anik and Bunt, 2021)), we map fairness perceptions to the model evaluation, deployment, and inference & usage stage. Most prior works have measured fairness perceptions after a (mock-up) model has been developed to ask stakeholders about certain outcomes or the AI system itself (Starke et al., 2022a). Beyond that, there are approaches trying to embed stakeholders’ values throughout conception and development of AI systems (e.g., (Noothigattu et al., 2018; Lee et al., 2019b; Stumpf et al., 2021; Yaghini et al., 2021)).
@@
The role of XAI. Explanations may serve stakeholders as a cue to confirm whether their ideas of fairness are implemented in a system. This idea has also been commonly expressed in prior literature (Ras et al., 2018; Shulner-Tal et al., 2022a, b; Papenmeier et al., 2019). For example, Lee et al. (2019a) indicate that explanations can decrease fairness perceptions when they reveal information that stands in conflict with peoples’ fairness beliefs. However, the effect of XAI on fairness perceptions is highly context-dependent and moderated by human factors like political ideology and self-interest (Starke et al., 2022b). This indicates that stakeholders require tailored information addressing their case-specific concerns. Affected parties, e.g., seem to appreciate information beyond a model’s inner workings such as system context, usage, and data (Schmude et al., 2023). We note that optimizing XAI to stimulate positive fairness perceptions isolated from complementary desiderata (e.g., trustworthiness or formal fairness) can lead to undesirable effects such as placebic explanations (Eiband et al., 2019), fairwashing (Aïvodji et al., 2019) or deception (Le Merrer and Trédan, 2020; John-Mathews, 2022).
3.5. Fairness With Human Oversight
@@
“Research suggests that neither humans nor ML models are likely to achieve fairness working alone. Instead, human–ML augmentation, where humans and technology work together to perform organizational tasks jointly, is the most promising path to achieving fairness.” (Teodorescu et al., 2021, p. 2)
@@
Beyond the (formal) fairness of a model itself, fairness can also refer to the decision-making process the model is embedded in. Fairness with human oversight aims at installing and supporting a human decision-maker to realize case and context-sensitive fairness objectives through human oversight or human discretion.
@@
The human-AI setting can take various forms but usually involves overseeing and overruling unfair outputs or fostering effective reliance behavior (Teodorescu et al., 2021). Although legal and ethical guidelines often demand human oversight (European Commission. Directorate General for Communications Networks, Content and Technology. and High Level Expert Group on Artificial Intelligence., 2019; Enqvist, 2023), the effect of human oversight on fairness is not necessarily beneficial and not well understood yet (Walmsley, 2021; Langer, 2023; Teodorescu et al., 2021; Schoeffer et al., 2024). Because full automation is currently not permitted in many contexts, this desideratum does not necessarily capture fairness through human oversight, but potentially also fairness despite human oversight.
@@
Stage of the Lifecycle. Fairness with human oversight becomes relevant after the model has been deployed, i.e., during inference and usage.
@@
The Role of XAI. Where formal implementation of fairness during model development is difficult or human oversight is required, XAI is crucial to inform human discretion such that fairness objectives can still be realized. In this sense, XAI is commonly proposed to support human decision-makers and domain experts in fostering fairer decisions (e.g., (Stumpf et al., 2021; Slack et al., 2020; Nakao et al., 2022)). For example, if problematic behavioral patterns of users are observed, XAI could provide tailored information to reconsider decisions, similar to Miller’s (Miller, 2023) idea of “evaluative AI”. Beyond simple recommendations, such information may include comparison to similar instances (Chakraborty et al., 2020), disclosure of uncertainty (Bhatt et al., 2021), or conditional heatmaps in computer vision tasks (Achtibat et al., 2023). However, both designing XAI and training human decision-makers to interact with the explanations is challenging (Lebovitz et al., 2022). As feature importance explanations may even hinder fairness of human decisions (Schoeffer et al., 2024), we are in need of more conceptual and empirical research on how to design XAI towards fairness objectives in human-in-the-loop settings (e.g., how to effectively override certain types of AI recommendations that violate fairness objectives).
3.6. Empowering Fairness

### v4

@@ 3 How XAI Can Be Leveraged for Fairness Along the AI Lifecycle
Grounded in a broad interdisciplinary literature review (summarized in Table 1), we distill eight categories of what previous work on fairness has called for. We call these categories fairness desiderata, and each desideratum can be instantiated with specific fairness objectives. For the purposes of this paper, we focus specifically on fairness desiderata connected to XAI. While various stakeholders are involved in utilizing XAI for fairness, an extensive discussion of their various roles is beyond the scope of this paper. When we mention stakeholders, we rely on the taxonomy provided by Langer et al. [40].
@@
Table 1: Fairness desiderata and their related concepts in existing interdisciplinary literature.
@@
Understanding bias [2]
Interdisciplinary fairness conceptualization [12]
Lessons from political philosophy [19]
Socio-technical perspective [64]
@@
Sampling bias [65, 3]
Data errors and bias [66]
Data-centric factors in algorithmic fairness [67]
@@
Algorithmic bias [3]
Formal fairness definitions [68]
Fairness metrics [15]
Statistical fairness criteria [13]
Disparate impact & disparate treatment [1]
@@
Fairness perceptions [4, 69]
Perceptions of justice [70, 71]
Fairness judgment [72]
Society-in-the loop [73]
@@
Human-ML augmentation for fairness [74]
Appropriate reliance [75]
Human-in-the-loop [72]
Domain expertise [76]
@@
Self-informed advocacy [77]
Procedural fairness [78]
Counterfactual explanations [79]
Fair and adequate explanations [80]
The explanation game [81]
@@
Long-term effects of algorithmic fairness [3]
Fairness drift [82]
Fairness through time [83]
Fairness monitoring [84, 85]
@@
Informational fairness [70]
Design publicity [52]
Model cards [49]
Outward transparency [86]
Figure 2: Fairness desiderata along the AI lifecycle depicting how XAI may directly contribute to these desiderata and how it may contribute to informational fairness across all desiderata (symbolized by speakerphones).
3.1 Fairness Understanding
“Accounting for bias not only requires an understanding of the different sources, that is, data, knowledge bases, and algorithms, but more importantly, it demands the interpretation and description of the meaning, potential side effects, provenance, and context of bias.” [2, p. 8]
@@ 3.1 Fairness Understanding
Among others, this may include a sound understanding of fairness schools of thought [12], existing social inequalities [53], stakeholder-centered fairness requirements [40], or relevant legal frameworks [87]. As the following fairness desiderata will show, there is no one-size-fits-all way in which a system can be fair, both across (e.g., a system can be perceived as fair despite not being formally fair) as well as within fairness desiderata (e.g., formal fairness criteria pose inherent tradeoffs [27]).
@@
The Role of XAI. XAI can contribute to a better understanding of fairness across all fairness desiderata which, e.g., may lead to a re-evaluation of earlier fairness objectives. Data-centric explanations [88] might reveal pre-existing group disparities (e.g., in the form of differing base rates such as the gender pay gap), which are crucial factors in determining the fairness objectives and how to achieve them. Further, simulating and testing various formal fairness metrics promotes an understanding of conflicting objectives that may lead to adjustments or re-weightings of fairness objectives [3]. Similarly, XAI may change stakeholders’ view of proactively using sensitive features in a specific decision-making context [62]. Concerning “substantive” fairness [53, 21], XAI may also be applied to normatively gauge the legitimacy of certain features depending on the societal context.
3.2 Data Fairness
@@
“Explanations […] are crucial for helping system developers and ML practitioners to debug ML algorithms by identifying data errors and bias in training data, such as measurement errors and misclassifications, data imbalance, missing data and selection bias, covariate shift, technical biases introduced during data preparation, and poisonous data points injected through adversarial attacks.” [66, p. 248]
@@
Undesirable model behavior often stems from flawed data that contains misrepresentations of the world (e.g., erroneous, mislabeled, or imbalanced data) or accurate representations that are societally undesirable (e.g., historical inequalities). The goal of data fairness is to identify and address such flaws in the data set used to train an AI model.
@@
Prior studies have pointed out data issues as drivers of unfairness [89, 65, 67] or highlighted data as a starting point for unfairness mitigation [90, 91, 92]. Mehrabi et al. [65] provide a comprehensive list of data biases that may introduce unfairness early on in AI development. Awareness and understanding of these biases are crucial to derive strategies to handle them.
@@
Stage of the Lifecycle. Data fairness relates to data collection (e.g., in the form of labeling errors, sampling bias, imbalances, etc.) and data analysis, which aims to identify and potentially mitigate unfair data characteristics early on. However, data fairness also reaches into feature selection which is usually part of an iterative loop together with model construction. For example, features might be dropped if they are not justifiably task-relevant. Importantly, in the context of sensitive attributes, developers should always be aware of the flaws of the idea of “fairness through unawareness” [16, 62].
@@
The Role of XAI. As data is a main source of unfairness [65], XAI is suited to identify potential disparities, imbalances, or abnormalities manifested in the available data early on. Descriptive statistics are a natural first step to explore pre-existing disparities [67]. Anik and Bunt [88] and Mitchell et al. [49] provide two examples of how to present simple descriptions and visualizations about the collection, feature distributions, and patterns of a dataset. XAI techniques have further been claimed to reveal instances and features in the data that have undesirable effects on the model output [93, 94, 66] which, however, are often subject to questionable causal and normative assumptions [95, 96]. These approaches indicate a strong connection between data fairness and formal fairness.
3.3 Formal Fairness
@@
“To counteract biases, it is, therefore, crucial to enable their detection. Explainability approaches may aid in this regard by providing means to track down factors that may have contributed to unfair and unethical decision-making processes and either to eliminate such factors, to mitigate them, or at least to be aware of them.” [40, p. 6]
@@
The most common fairness desideratum is concerned with formal model properties. By formal fairness, we refer to the vast array of formal criteria that have been proposed as mathematical and statistical measures of fairness [13, 15].
@@
Consistent with Verma and Rubin [14], this includes all fairness definitions based on statistical measures (e.g., demographic parity), similarity measures (e.g., fairness through awareness), or based on causal reasoning (e.g., counterfactual fairness). Formal fairness notions are often distinguished into group and individual fairness. Group fairness criteria typically require a form of parity between demographic groups, e.g., along sensitive attributes like gender or race [54]. Individual fairness criteria typically demand to treat similar people alike [16].
@@
Stage of the Lifecycle. Formal fairness is particularly relevant for the iterative loop of model construction and model evaluation. As soon as the first model prototype is ready, fairness metrics can be evaluated. Based on the evaluation, unfairness mitigation techniques can be implemented and validated iteratively [97].
@@
The Role of XAI. Regarding formal fairness, XAI could be of exploratory value, offering a plethora of tools and a novel perspective from which to explore formal fairness from multiple angles. Again, descriptive statistics present a natural first step to identify potential disparities, imbalances, or abnormalities outputted by the model [54, 98]. This is especially useful if formal fairness objectives are already specified, e.g., when testing whether formal group fairness metrics are sufficiently satisfied [97]. XAI might also point towards specific features driving the violation of group fairness metrics [99] or shed light on more subtle forms of formal fairness such as fairness of recourse [100]. XAI has further been claimed to elucidate the complex interplay between “task-relevant” and correlated “protected” features (e.g., [101]) which, again, relies on causal and normative assumptions [95, 96]. Technical possibilities of XAI to explore formal fairness are manifold but require utmost caution when applied for unfairness mitigation in specific application contexts and should not be misinterpreted as fairness “proofs” or “guarantees” [6]. Particularly, when interpreting the legitimacy of using sensitive information, it is paramount to account for the “fairness through unawareness” fallacy [16, 61] and for differential subgroup validity [102, 103].
3.4 Perceived Fairness
@@
“We argue that only fair systems that are also perceived to be fair by their users should be accepted and employed in practice.” [104, p. 5]
@@
Another common fairness desideratum refers to positive perceptions of stakeholders, which is often related to trust and acceptance [104, 105]. In this sense, perceived fairness captures how stakeholders, particularly those affected by a decision, perceive the fairness of the AI system.
@@
Measures of perceived fairness can be derived from the justice constructs of Colquitt [70], which decompose fairness perceptions into a procedural, distributive, and informational dimension (see also [71, 72, 69]). Accounting for fairness perceptions promotes a valuable means to design AI systems based on human needs and ideals while giving voice to societal values, which Rahwan [73] coined as “society-in-the-loop”. Notably, the desirable aspect of perceived fairness does not necessarily or solely lie in positive fairness perceptions but in “appropriate” fairness perceptions [106].
@@
Stage of the Lifecycle. Although fairness perceptions can be relevant at any stage of the lifecycle (e.g., to evaluate data fairness [88]), we map fairness perceptions to the model evaluation, deployment, and inference & usage stage. Most prior works have measured fairness perceptions after a (mock-up) model has been developed to ask stakeholders about certain outcomes or the AI system itself [4]. Beyond that, there are approaches trying to embed stakeholders’ values throughout the conception and development of AI systems (e.g., [107, 108, 109, 110]).
@@
The Role of XAI. Explanations may serve stakeholders as a cue to confirm whether their ideas of fairness are implemented in a system. This idea has also been commonly expressed in prior literature [111, 104, 112, 105]. For example, Lee et al. [113] indicate that explanations can decrease fairness perceptions when they reveal information that stands in conflict with peoples’ fairness beliefs. However, the effect of XAI on fairness perceptions is highly context-dependent and moderated by human factors like political ideology and self-interest [114]. This indicates that stakeholders require tailored information addressing their case-specific concerns. Affected parties, e.g., seem to appreciate information beyond a model’s inner workings, such as system context, usage, and data [115]. We note that optimizing XAI to stimulate positive fairness perceptions isolated from complementary desiderata (e.g., trustworthiness or formal fairness) can lead to undesirable effects such as placebic explanations [116], fairwashing [56] or deception [117, 118].
3.5 Fairness With Human Oversight
@@
“Research suggests that neither humans nor ML models are likely to achieve fairness working alone. Instead, human–ML augmentation, where humans and technology work together to perform organizational tasks jointly, is the most promising path to achieving fairness.” [74, p. 2]
@@
Beyond the (formal) fairness of a model itself, fairness can also refer to the decision-making process in which the model is embedded. Fairness with human oversight aims at installing and supporting a human decision-maker to realize case and context-sensitive fairness objectives through human oversight or human discretion.
@@
The human-AI setting can take various forms but usually involves overseeing and overruling unfair outputs or fostering effective reliance behavior [74]. Although legal and ethical guidelines often demand human oversight [78, 119], the effect of human oversight on fairness is not necessarily beneficial and not well understood yet [86, 120, 74, 43]. Accordingly, this desideratum does not necessarily capture fairness through human oversight, but potentially also fairness despite human oversight.
@@
Stage of the Lifecycle. Fairness with human oversight becomes relevant after the model has been deployed, i.e., during inference & usage.
@@
The Role of XAI. Where formal implementation of fairness during model development is difficult or human oversight is required, XAI is crucial to inform human discretion so that fairness objectives can still be realized. In this sense, XAI is commonly proposed to support human decision-makers and domain experts in fostering fairer decisions (e.g., [109, 121, 122]). Beyond simple recommendations, such information may include a comparison to similar instances [76], disclosure of uncertainty [123], or conditional heatmaps in computer vision tasks [124]. To tackle automation bias, Miller [125]’s concept of “evaluative AI” also highlights providing not only explanations for a certain recommendation but rather balanced evidence for multiple possible outcomes. However, both designing XAI and training human decision-makers to interact with the explanations is challenging [126]. As feature importance explanations may even hinder fairness of human decisions [43], we are in need of more conceptual and empirical research on how to design XAI towards fairness objectives in human-in-the-loop settings (e.g., how to effectively override certain types of AI recommendations that violate fairness objectives).
3.6 Empowering Fairness